Consider a box with two chambers labeled 1 and 2. Each chamber is filled with identical ideal gases that are initially at the same pressure,

temperature, and volume; and the vertical wall between the chambers is free to move without transferring any thermal energy or gas between

the chambers. Each chamber starts at an initial volume of 2 m^3 and pressure 100 kPa. We then heat the gas of Chamber 1, which causes the

dividing wall to move. When the wall comes to rest, Chamber 1 has an absolute temperature that is 50% higher than Chamber 2. Given that

the dividing wall has a cross-sectional area of 1 m^2, how far has the wall moved from its original position?
Answer #1

Solution Let the partition wall moved by X m So New volume of Chamber I am v=2m3 + xm3 = (2+xm- and new volume at Chamber2 =